Eduardo created a metal containing 35% gold by combining two other metals. One of these other metals weighed 2 kg and contained 15% gold. If the other weighed 8kg then what percent of it was gold?
I believe I would set up an equation .35(x + .15) = 2(.15) +8x where x is % of gold in metal 2, but I don't know if that's right, or where to go from there

Answer by stanbon(75887)

Eduardo created a metal containing 35% gold by combining two other metals. One of these other metals weighed 2 kg and contained 15% gold. If the other weighed 8kg then what percent4 of it was gold?
-----
Equation:
gold + gold = gold
0.15*2 + x(8) = 0.35(2+8)
-------
Multiply thru by 100 to get:
15*2 + 100x*8 = 35*2 + 35*8
---
30 + 800x = 70 + 280
800x = 320
x = 32/80 = 4/10 = 40%
---
Cheers,
stan H.
==============